OpenStudy (kearsten):

medal and fan solve for x x^2=-9/4

OpenStudy (lfreeman1285):

because we're trying to get x all by itself, we need to square root both sides. \[\sqrt{x}=\sqrt{-9/4}\] This can be simplified to\[\sqrt{-1 * 3 *3}\div \sqrt{4}\] The square root of 4 can be easily simplified to 2, and the pair of threes can be taken out from under the square root. Now you have \[3\sqrt{-1}\div2\] There's a variable for the square root of negative one, known as i. We use i because negative numbers aren't able to be square rooted. Therefore, your answer is 3i/2

OpenStudy (lfreeman1285):

Sorry if that's a little confusing, I'd be happy to clarify anything for you!

OpenStudy (anonymous):

@hype.child @whatnow @JadedInsomniac @LeftFielding @Learner11 @LifeIsADangerousGame @Hero @GIL.ojei 1 one ta answer da questions ill medal 1